Can someone please correct my answer or answer the question? Please and thank you :) At a recent seminar you attended, the invited speaker was discussing some of the advantages and disadvantages of standard cosComplete this question by entering your answers in the tabs below. Required 1 Required 2 Prepare a table that contains the stComplete this question by entering your answers in the tabs below. Required 1 Required 2 Compute the direct labor efficiency
Answer #1

2.

Direct labor efficiency variance = (Standard hours - Actual hours) * Standard rate

Direct labor efficiency variance = (5,800*3.26536 - 24,500) * $29

Direct labor efficiency variance = (18,939 - 24,500) * $29 = $161,269 Unfavorable

Answer provided to Requirement 1 is correct.